OpenStudy (anonymous):

If f(x) = 1 over 2x + 2 and g(x) = −3x − 1, find g[f(4)].

OpenStudy (anonymous):

put x=4 on 1/2x+2

OpenStudy (anonymous):

and give that answer as value of x in g(x)

OpenStudy (anonymous):

i got 10 that doesnt seem right

OpenStudy (anonymous):

\[f(x) = \frac{1}{2x + 2} \] & g(x) = −3x − 1 \[f(x) = \frac{1}{2x + 2} \rightarrow f(4) = \frac{1}{2\times4 + 2} =\frac{1}{10}\] \[ f(4) = \frac{1}{10}\] Now g(x) = −3x − 1 substituting x= f(4) in g(x) we find \[g[f(4)] =-3f(4)-1= -3\times \frac{1}{10} -1=\frac{-3}{10} -1=\frac{-3-10}{10}=\frac{-13}{10}\]